Perez’s Doll Company produces handmade dolls. The standard amount of time spent on each doll is 1.50 hours. The standard cost of labor is $7.73 per hour. The company planned to make 8,400 dolls during the year but actually used 13,100 hours of labor to make 9,900 dolls. The payroll amounted to $104,800.

Required

  1. Should labor variances be based on the planned volume of 8,400 dolls or the actual volume of 9,900 dolls?

  2. Prepare a table that shows the standard labor price, the actual labor price, the standard labor hours, and the actual labor hours.

  3. Compute the labor price variance and indicate whether it is favorable (F) or unfavorable (U).

  4. Compute the labor usage variance and indicate whether it is favorable (F) or unfavorable (U).
Answer #1

Solution 1:

Labor variances should be based on the actual volume of 9,900 dolls.

Solution 2:

Labor Variance Information Table
Standard Labor Price $7.73 Per Hour
Actual Labor Price ($104,800/13100) $8.00 Per Hour
Standard labor Hours (9900*1.5) 14850 Hours
Actual Labor Hours used 13100 Hours

Solution 3:

Labor Price variance = (SP- AP) * AH = ($7.73 - $8) * 13100 = $3,537 Unfavorable

Solution 4:

Labor Usage variance = (SH - AH) * SR = (14850 - 13100) * $7.73 = $13,527.50 Favorable